10) Ethyl iodide (C2H5I) decomposes at a certain temperature in the gas phase as follows:

C2H5I(g) → C2H4(g) + HI(g)

From the following data determine the order of the reaction and the rate constant.

Time (min) [C2H5I] (M)

0.0 2.00

15.0 1.82

30.0 1.64

48.0 1.42

75.0 1.10

1) Reaction order:

a) 0

b) 1

c) 2

2) Rate Constant: (enter only the numerical value of the rate constant). _______
Answer #1

first 15 minutes it was consumed =2-1.82   = 0.18M

rate = change in concentration/time taken

         = 0.18/15   = 0.012M/min

next 30 to 48 minutes ( 18 minutes)

Rate = change in concentration/time taken

           = 1.64-1.42/18

           = 0.22/18   = 0.012M/min

from 48 to 75 ( 75-48 = 27min)

Rate = change in concentration/time taken

         =   1.42-1.1/27

          = 0.32/27   = 0.012M/min

The rate of reaction is same. so the rate of reaction does not depends upon the concetration of reaction.It is zero order reaction

The reaction order is zero

a.0 >>>>answer

Rate = k[C2H5I]0

Rate = K

K   = 0.012M/min >>>>answer